OpenStudy (anonymous):

Express answers in simplest exact form. What is the area of a sector with radius equal to 8 and measure of arc equal to 300°? A = ? PI

OpenStudy (anonymous):

(5pi/6)*64

OpenStudy (anonymous):

or 160pi/3

OpenStudy (amistre64):

5/6 the area of a circle with a radius of 8

OpenStudy (anonymous):

\[\frac{300}{360}\pi 8^2=\frac{160 \pi }{3}=167.552 \]